the coordinates of point P and Q are (-2,3) and (1,2) respectively .A straight line QR is perpendicular ti the straight line PQ.find the equation of the straight line QR
1
Expert's answer
2020-11-04T16:38:36-0500

"a(x-x_0)+b(y-y_0)=0\text{ where } \vec n\text{ normal line vector}"

"(x_0,y_0)=Q(1,2)"

"\vec n=\vec {PQ}=(1-(-2),2-3)=(3,-1)"

"3(x-1)-1*(y-2)=0"

"3x-y-1=0"

Answer: "3x-y-1=0"